### v2

# Style Content Decomposition-based Data Augmentation for Domain Generalizable Medical Image Segmentation

Abstract
Due to domain shifts across diverse medical imaging modalities, learned segmentation models often suffer significant performance degradation during deployment. These domain shifts, typically caused by variations in imaging systems, generally comprise two principal components: 1) ”style” shifts, referring to global disparities in image properties such as illumination, contrast, and color; and 2) ”content” shifts, which involve local discrepancies in anatomical structures. To address domain shifts in medical image segmentation, a core challenge arises: how can we decouple the factors within images that determine their ”style” and ”content” components? To this end, we first propose a linear style-content decomposition method that factorizes an image into style codes and content maps, explicitly modeling the ”style” and ”content” components. Building on this, we introduce a Style-Content decomposition-based data augmentation algorithm (StyCona), which leverages this decomposition strategy to guide augmentation of both the global style and local content of source-domain images, enabling the training of a well-generalized model for domain-generalizable medical image segmentation. StyCona is a simple yet effective plug-and-play module that substantially improves model generalization without requiring additional training parameters or modifications to segmentation model architectures. Experiments on cardiac magnetic resonance imaging and fundus photography segmentation tasks, with single and multiple target domains respectively, demonstrate the effectiveness of StyCona and its superiority over state-of-the-art domain generalization methods. The code will be released at https://github.com/Senyh/StyCona.
Introduction
Medical image segmentation is critical for computer-aided diagnosis. Driven by large and diverse annotated datasets, deep learning-based segmentation models have achieved remarkable progress (Ronneberger, Fischer, and Brox 2015; Milletari, Navab, and Ahmadi 2016). However, data distribution mismatches (a.k.a., domain shifts), caused by diverse imaging protocols, equipment vendors, or patient populations, etc., between source (training) and target (testing) domains hinder the generalizability of trained models for clinical deployment (Castro, Walker, and Glocker 2020; Zhou et al. 2022; Guan and Liu 2021).
Figure 1: Illustration of style codes and content maps for bSSFP and LGE MRI sequences: a) original images showing disparities in both style (global image appearance) and content (local anatomical structures), b) image histograms reflecting domain shifts in terms of pixel intensities, c) style-swapped images generated by swapping the two original images’ style codes (top: bSSFP image with LGE style; bottom: LGE image with bSSFP style), d) content maps of the original images, and e) t-SNE (Van der Maaten and Hinton 2008) visualization of domain shifts (top: style shift; bottom: content shift).
Domain generalization (DG) has emerged as a promising strategy to mitigate domain shifts by training models solely on source-domain data and generalizing them to unseen domains. Current domain generalization (DG) methods encourage segmentation models to learn domain-invariant representations through representation learning (Chen et al. 2019, 2020; Pei et al. 2021; Peng et al. 2022; Gao et al. 2022, 2023), data augmentation (a.k.a., domain randomization) (Ouyang et al. 2022; Zhou et al. 2021; Chen et al. 2023, 2025), as well as the combination of both schemes (Gu et al. 2023; Yang and Soatto 2020; Xu et al. 2021a; Li et al. 2023). Specifically, the first line of work focuses on explicitly learning domain-invariant information via deterministic (Chen et al. 2019, 2020; Pei et al. 2021) or statistical (Gao et al. 2022, 2023) modeling. While these methods yield intuitive decoupled features on source-domain data, domain shifts pose a dilemma: the disentanglement models, constrained by domain-specific training, often fail to generalize their disentanglement capabilities across different domains. In contrast, data augmentation-based DG methods can circumvent this issue by expanding the source-domain data distribution and thus implicitly encouraging models to capture domain-invariant information. These methods leverage techniques such as random convolution (Xu et al. 2021b; Ouyang et al. 2022; Choi et al. 2023), and feature statistics editing (Zhou et al. 2021; Chen et al. 2023, 2025), etc. The third line of research employs learning-based (Gu et al. 2023) or rule-based (Yang and Soatto 2020; Xu et al. 2021a; Li et al. 2023) disentanglement to decompose images into domain-invariant contents and domain-specific styles, followed by style augmentation based on the disentangled components. However, the former depends on additional encoders and decoders, as well as access to multiple domain images and their corresponding labels. These requirements are computationally intensive and may be infeasible in practice, e.g., single-source domain generalization or domain-label unavailable scenarios. In contrast, the latter assumes invariance in the Fourier phase spectrum, which limits its effectiveness for data augmentation to alleviate domain shifts. Moreover, the specific components contributing to domain shifts remain underexplored. This raises a question: Can a deeper understanding of the underlying nature of domain shifts guide the design of more effective data augmentation methods for domain generalizable medical image segmentation?
To answer this question, we begin with a visual investigation of domain shifts in medical images. As shown in Fig.1(a-b), domain shifts generally manifest as variations in pixel intensity, comprising two major components: 1) style shift, referring to global disparities in image characteristics such as brightness, contrast, and color; and 2) content shift, referring to local anatomical discrepancies caused by differences in tissue visualization across imaging systems. Inspired by these observations, rather than learning-based (Gu et al. 2023) or rule-based (Yang and Soatto 2020; Xu et al. 2021a; Li et al. 2023) disentanglement, we explicitly model the ”style” and ”content” components of an image by linearly decomposing it into a style code and content maps. We reveal that 1) the style code captures global image characteristics within a given anatomical structure and 2) the content maps describe the image’s anatomical structure. To provide a clearer illustration, we visualize style-swapped images (by exchanging two images’ style codes) and content maps in Fig. 1(c-d); it can be observed that the content maps delineate the anatomical structures of the original bSSFP and LGE images, and the style-swapped images transfer appearance between each other while preserving the original anatomical structures. Furthermore, the shift observed in the style codes is more significant than that in the content maps [Fig. 1(e)], as global variations in image characteristics are generally more substantial than local differences in anatomical structures. These results align with the aforementioned intuition. Based on this, we categorize domain shifts into two components: 1) style shifts (global image property variations) as indicated by deviations in the style codes, and 2) content shifts (local anatomical structure discrepancies) as reflected in differences between the content maps. Given that explicitly measuring and reducing these two types of shifts is infeasible due to the unavailability of target domain data during training, we instead aim to perturb the ”style” and ”content” components of source domain images for generating images from diverse domains (simulating patients undergoing different imaging systems) to train a domain generalizable medical image segmentation model.
To this end, we propose a style content decomposition-based data augmentation algorithm (StyCona) to advance domain generalizable medical image segmentation. Specifically, StyCona performs perturbations on the ”style” (global image characteristics) and ”content” (local anatomical structures) by 1) blending style codes and 2) mixing content maps. We evaluate StyCona on cross-domain cardiac magnetic resonance imaging (MRI) segmentation and optic cup (OC)/optic disk (OD) fundus photography segmentation tasks. Experimental results demonstrate that StyCona is a promising solution for domain generalizable medical image segmentation and achieves state-of-the-art performance.
Our main contributions can be summarized as follows:
  • $\bullet$
    To the best of our knowledge, this is the first work that decomposes domain shifts into style shifts and content shifts through linear decomposition of images in domain generalizable medical image segmentation tasks.
  • $\bullet$
    We propose StyCona, a novel data augmentation algorithm for domain generalizable medical image segmentation. Our method can be easily integrated into off-the-shelf medical image segmentation backbones and effectively mitigates the two types of domain shifts.
  • $\bullet$
    StyCona leads to state-of-the-art performance across a wide range of tasks and provides a compelling alternative to Fourier transformation-based, random convolution-based, and feature statistics editing-based domain generalization methods.
Related work
Domain Generalization
Domain generalization (DG) aims to train a model on a source domain and enable it to generalize to unseen target domains. Numerous DG methods have been proposed for both natural and medical image scenarios (Zhou et al. 2022; Castro, Walker, and Glocker 2020). Specifically, these methods enforce the learning of domain-invariant representations through various approaches, including: deterministic or statistical modeling of domain-invariant and domain-specific features (Gu et al. 2023; Gao et al. 2022, 2023); frequency transformation-based mix-up, which combines Fourier amplitude spectra from different images for style augmentation (Yang and Soatto 2020; Xu et al. 2021a; Li et al. 2023; Zhao et al. 2024); random convolution-based data augmentation, which perturbs pixel intensities via convolution neural networks with random parameters (Xu et al. 2021b; Ouyang et al. 2022; Choi et al. 2023); as well as feature statistics editing-based style transfer, which perturbs image- or feature-level statistics (e.g., mean and standard deviation) to achieve style or content augmentation (Zhou et al. 2021; Chen et al. 2023, 2025). These techniques are generally applicable to both image recognition and semantic segmentation tasks, with the latter requiring additional consideration for transformation equivalence. However, the fundamental nature of domain shifts, crucial for guiding the development of more effective DG approaches, remains underexplored. In this paper, we propose a novel framework that first decomposes domain shifts into style and content shifts through linear modeling of images and then designs a novel data augmentation algorithm to mitigate both types of domain shifts.
Content-Style Decomposition
Content-style decomposition disentangles features into domain-invariant content and domain-specific style components for addressing the domain shift issue. In the context of semantic segmentation, this technique has been widely employed in the design of both unsupervised domain adaptation (Chen et al. 2019, 2020; Pei et al. 2021; Peng et al. 2022; Yang and Soatto 2020) and domain generalization (Gu et al. 2023; Gao et al. 2022, 2023; Xu et al. 2021a; Li et al. 2023; Zhao et al. 2024) frameworks. Research in this area generally falls into two categories: 1) learning-based (Chen et al. 2019, 2020; Pei et al. 2021; Peng et al. 2022; Gu et al. 2023; Gao et al. 2022, 2023) and 2) rule-based approaches (Yang and Soatto 2020; Xu et al. 2021a; Li et al. 2023; Zhao et al. 2024). Learning-based methods rely on additional deep neural network-based encoders and decoders, with specific constraints (e.g., contrastive learning on decoupled components (Gu et al. 2023)), to achieve feature disentanglement either deterministically (Chen et al. 2019, 2020; Pei et al. 2021; Peng et al. 2022; Gu et al. 2023) or statistically (Gao et al. 2022, 2023). Rule-based methods adopt the Fourier transformation to perform direct decomposition. However, they impose the constraint that phase spectra should be fixed to ensure the invariance of the original contents during reconstruction. This constraint limits their flexibility and effectiveness for data augmentation. In contrast, we propose a simple yet effective content-style decomposition method that linearly factorizes an image into a style code and content maps, enabling style-content augmentation through perturbations on the decoupled components.
Figure 2: Schematic diagram of the proposed style content decomposition-based data augmentation algorithm (StyCona). StyCona includes three steps: 1) singular value decomposition, 2) style augmentation, and 3) content augmentation. $x_{i}$ and $x_{j}$ represent original images, and $\bar{\tilde{x}}$ denotes a StyCona-augmented sample.
Methodology
For a domain generalizable medical image segmentation task, the training set can be denoted as $\mathcal{D}=\{\mathcal{S}_{k}|\mathcal{S}_{k}=\{(x_{i(k)},y_{i(k)})_{i=1}^{N_{k}}\}_{k=1}^{K}\}$, where $x_{i(k)}$ represents the $i^{th}$ image in the ${k^{th}}$ domain and $y_{i(k)}$ is its corresponding ground truth segmentation label. This paper focuses on the most challenging single-source domain generalization setting (i.e., $K=1$), in which the segmentation model $f(\cdot;\theta)$ is trained on a single source domain and evaluated on one or more unseen target domains. Note that we omit the domain index for brevity.
Overview
We assume that domain shifts in medical images resulting from variations in image acquisition processes generally comprise two principal components: 1) ”style” shifts, referring to global variations in image properties such as illumination, contrast, and color; and 2) ”content” shifts, which involve local changes in image characteristics and are usually associated with differences in the visibility of specific tissues. Under this assumption, a key problem arises for addressing the domain shifts: how can we decouple the factors within images that determine the ”style” and ”content” components? Although several approaches have been proposed to address this problem, they either rely on auxiliary disentanglement networks (Gu et al. 2023; Gao et al. 2022, 2023) or impose constraints on the perturbation of the decoupled components (Xu et al. 2021a). Motivated by the effectiveness of linear decomposition in image compression (Andrews and Patterson 2003; Ranade, Mahabalarao, and Kale 2007) and restoration (Zhu et al. 2014; Guo et al. 2015; Gao and Zhuang 2020), we propose to linearly decompose an image into its style code and content maps, from which we validate both aforementioned types of domain shifts, as shown in Fig. 1. Based on the linear decomposition, we introduce a style content decomposition-based data augmentation algorithm, termed StyCona, for domain generalizable medical image segmentation. The training process for leveraging StyCona to train a domain-generalizable medical image segmentation model is summarized in Algorithm 1.
Style Content Decomposition-based Data Augmentation (StyCona)
The core idea behind StyCona is that the style and content shifts can be mitigated through style and content augmentations, respectively. To this end, the StyCona algorithm mainly includes two components: 1) image decomposition and 2) style content augmentation. Specifically, the image decomposition is achieved using singular value decomposition (SVD) (Klema and Laub 1980) to factorize an image into its singular values (style codes) and rank-one matrices (content maps) The style content augmentation is then performed by perturbing style codes and content maps, respectively. As illustrated in Fig. 2, we realize the StyCona algorithm with three steps: 1) singular value decomposition, 2) style augmentation, and 3) content augmentation.
Singular Value Decomposition
For an image $x$, its SVD is formulated as:
$x=\sum_{k}\underbrace{\sigma^{k}}_{style\,code}\,\,\underbrace{\boldsymbol{u}^{k}{\boldsymbol{v}^{k}}^{\mathsf{T}}}_{content\,map}$ (1)
where each singular value $\sigma^{k}$ acts as a style code and the rank-one matrix $\boldsymbol{u}^{k}{\boldsymbol{v}^{k}}^{\mathsf{T}}$ serves as a content map for the subsequent style and content augmentation.
Algorithm 1 StyCona for domain generalizable medical image segmentation
Input: $\mathcal{D}=\{(x_{i},y_{i})_{i=1}^{N}\}$
Parameter: $\theta$ for the segmentation model $f(\cdot;\theta)$
Output: Trained model $f(\cdot;\theta)$
1:for each iteration do
2:  # 1) Extract style codes and content maps from current batch original image $x_{i}$ and auxiliary image $x_{j}$ ($i\neq j$) by Eq. (1)
3:  $\boldsymbol{\sigma}_{i}\leftarrow x_{i}\,\,and\,\,\{\boldsymbol{u}^{r}_{i}\times\boldsymbol{v}^{r}_{i}\}\leftarrow x_{i}$
4:  $\boldsymbol{\sigma}_{j}\leftarrow x_{j}\,\,and\,\,\{\boldsymbol{u}^{r}_{j}\times\boldsymbol{v}^{r}_{j}\}\leftarrow x_{i}$
5:  # 2) Style Augmentation by Eq. (2)
6:  $\tilde{\sigma}_{i}^{k}=\alpha\times\sigma_{i}^{k}+(1-\alpha)\times\sigma_{j}^{k}$
7:  # 3) Content Augmentation by Eq. (3)
8:  $\bar{\boldsymbol{u}}_{i}^{k}=\beta\times\boldsymbol{u}_{i}^{k}+(1-\beta)\times\boldsymbol{u}_{j}^{k}\quad or\quad\bar{\boldsymbol{v}}_{i}^{k}=\beta\times\boldsymbol{v}_{i}^{k}+(1-\beta)\times\boldsymbol{v}_{j}^{k}$
9:  # 4) Obtain StyCona images via inverse singular value decomposition
10:  $\bar{\tilde{x}}_{i}=\sum^{r_{i}}_{k=1}\tilde{\boldsymbol{\sigma}}_{i}^{k}\bar{\boldsymbol{u}}_{i}^{k}{\bar{\boldsymbol{v}}_{i}^{k}}{}^{\mathsf{T}}$
11:  # 5) Forward propagation to obtain the predictive label for the StyCona image $\bar{\tilde{x}}_{i}$
12:  $\hat{\bar{\tilde{y}}}_{i}\leftarrow f(\bar{\tilde{x}}_{i},\theta)$
13:  # 6) Loss supervision by Eq. (4)
14:  Back-Propagate $\mathcal{L}$
15:  Update $\theta$
16:  return $f(\cdot;\theta)$
17:end for
Style Augmentation
Style augmentation aims to alleviate style shifts (global variations in image properties). As depicted in Fig. 2, we perform style augmentation on an image $x_{i}$ by blending its style codes (i.e., singular values) with those of an auxiliary image $x_{j}$ in an element-wise manner:
$\tilde{\sigma}_{i}^{k}=\alpha\times\sigma_{i}^{k}+(1-\alpha)\times\sigma_{j}^{k}$ (2)
where the weight $\alpha\sim U(0,1)$ controls the mixing strength, $\sigma_{i}^{k}$ represents the $k_{th}$ singular value of image $x_{i}$, $i/j$ indicates the sample index and $i\neq j$. After that, a style-augmented image $\tilde{x}_{i}$ can be generated using the perturbed singular values and the original singular vectors by Eq. (1).
Content Augmentation
We devise image content augmentation to mitigate content shifts (local disparities in anatomical structure). This is achieved by mixing a set of $t$ randomly selected content maps (i.e., rank-one matrices) from an image $x_{i}$ with another set of $t$ randomly selected content maps from an auxiliary image $x_{j}$ [Fig. 2]. The perturbation to each content map is formulated as:
$\bar{\boldsymbol{u}}_{i}^{k}=\beta\times\boldsymbol{u}_{i}^{k}+(1-\beta)\times\boldsymbol{u}_{j}^{k}\quad or\quad\bar{\boldsymbol{v}}_{i}^{k}=\beta\times\boldsymbol{v}_{i}^{k}+(1-\beta)\times\boldsymbol{v}_{j}^{k}$ (3)
where the weight $\beta\sim U(0,1)$ and $\boldsymbol{u}_{i/j}^{k}$/$\boldsymbol{v}_{i/j}^{k}$ represents the left/right singular vector of image $x_{i/j}$. Using the perturbed content maps and the original style codes, we can obtain a content-augmented image $\bar{x}_{i}$ via Eq. (1).
In a nutshell, by leveraging Eq. (1)(2)(3), StyCona conducts style and content augmentation and produces an augmented image $\bar{\tilde{x}}_{i}=\sum^{r_{i}}_{k=1}\tilde{\boldsymbol{\sigma}}_{i}^{k}\bar{\boldsymbol{u}}_{i}^{k}{\bar{\boldsymbol{v}}_{i}^{k}}{}^{\mathsf{T}}$.
Loss Supervision
Afterwards, we formulate the segmentation loss $\mathcal{L}$ for training a domain generalizable model $f(\cdot;\theta)$ based on StyCona images:
$\mathcal{L}=\frac{1}{N}\sum^{N}_{i}\mathcal{L}_{seg}(f(\bar{\tilde{x}}_{i},\theta),y_{i})$ (4)
where $\mathcal{L}_{seg}$ denotes a segmentation criterion.
Experiments and Results
Cross-Domain Cardiac Magnetic Resonance Imaging (MRI) Segmentation
The MS-CMR Dataset (Single Source and Single Target Domain) dataset (Zhuang 2018) contains 45 subjects, each with bSSFP and LGE MRI sequences, along with ground truth annotations for the right ventricle (RV), left ventricle (LV), and myocardium (MYO). All images are normalized to the range [0,1][0,1][ 0 , 1 ] and resampled to a uniform resolution of $\times 1.0$ 1.0 mm. This experimental setting evaluates a segmentation model’s ability to generalize across MRI sequences for cardiac structure segmentation, specifically from LGE to bSSFP and vice versa. For each direction, one sequence (e.g., bSSFP) is treated as the source domain and is divided into training and validation sets with a 8:28:28 : 2 ratio, while another sequence (e.g., LGE) serves as the target domain and is used solely for testing.
Cross-Domain Optic Cup (OC) and Optic Disc (OD) Fundus Image Segmentation
The Fundus Image Benchmark (Single Source and Multiple Target Domains(Chen et al. 2023) is established based on the ORIGA (Zhang et al. 2010), Drishti-GS (Sivaswamy et al. 2014), REFUGE (Orlando et al. 2020), and RIGA (Almazroa et al. 2018) (BinRushed and Magrabia) datasets. This setting evaluates the model’s generalization capability for the joint OC and OD segmentation across multiple target domains. The ORIGA dataset is used as the source domain with the training and validation split following TriD (Chen et al. 2023); the other datasets (BinRushed, Drishti-GS, Magrabia, and REFUGE) serve as target domains for cross-domain evaluation.
Implementation Details.
We conducted the experiments using PyTorch (Paszke et al. 2019) on an NVIDIA A40 GPU with 48G GPU memory. All compared methods were optimized using an AdamW optimizer (Kingma and Ba 2015) with a fixed learning rate of $1e-4$ during the 100 training epochs for both the cadiac MRI and fundus image segmentation tasks. U-Net (Ronneberger, Fischer, and Brox 2015) was employed as the baseline segmentation model. The combination of cross-entropy and Dice loss (Milletari, Navab, and Ahmadi 2016) was used as the segmentation criterion. All images were resized to $\times 256$ 256 for both training and testing. During evaluation, the predicted labels were rescaled to their original resolutions. We set the number of perturbed content maps $t=16$ (please refer to the ablation study for further analysis).
Comparison with State of the Arts
We extensively compared our StyCona with state-of-the-art DG methods: 1) Fourier transformation-based: amplitude swap (AmpSwap) (Yang and Soatto 2020), amplitude mixup (AmpMix) (Xu et al. 2021a), and FMAug (Li et al. 2023)); 2) Feature statistics editing-based: MixStyle (Zhou et al. 2021), TriD (Chen et al. 2023), and ConStyX (Chen et al. 2025); 3) Random convolution-based: RandConv (Xu et al. 2021b), CIDA (Ouyang et al. 2022), and PRandConv (Choi et al. 2023). We re-implemented all the compared methods in a unified experimental setup for fair comparison. All the methods are built upon the same baseline segmentation model (U-Net). In general, StyCona sets the state-of-the-art in all three settings, both quantitatively [Table 1 and Table 2] and qualitatively [Fig. 3], showcasing its superiority in domain generalizable medical image segmentation.
Table 1: Comparison with state-of-the-art methods on the cross-domain cardiac MRI segmentation task (single target domain). The best and second-best results are highlighted in bold and underline, respectively.
Method bSSFP $\rightarrow$ LGE LGE $\rightarrow$ bSSFP
DSC $\uparrow$ ASD $\downarrow$ DSC $\uparrow$ ASD $\downarrow$
U-Net (Ronneberger, Fischer, and Brox 2015) 65.97 6.55 76.91 3.41
AmpSwap (Yang and Soatto 2020) 70.93 6.48 78.74 3.26
AmpMix (Xu et al. 2021a) 63.89 6.44 80.07 2.46
FMAug (Li et al. 2023) 72.46 13.67 80.54 2.57
MixStyle (Zhou et al. 2021) 69.26 10.91 80.81 2.69
TriD (Chen et al. 2023) 65.12 10.25 79.14 4.60
ConStyX (Chen et al. 2025) 68.63 7.89 81.30 2.85
RandConv (Xu et al. 2021b) 71.90 5.65 75.17 4.27
CIDA (Ouyang et al. 2022) 69.72 8.63 72.85 4.61
PRandConv (Choi et al. 2023) 56.61 8.75 75.57 3.38
StyCona (ours) 73.39 4.33 81.59 2.24
Table 2: Comparison with state-of-the-art methods on the fundus segmentation tasks with 10% labeled data (Source: ORIGA; Target: REFUGE, Drishti-GS, BinRushed, and Magrabia). The best and second-best results are highlighted in bold and underline, respectively.
Method BinRushed Drishti-GS Magrabia REFUGE Average
DSC (%) $\uparrow$ ASD $\downarrow$ DSC (%) $\uparrow$ ASD $\downarrow$ DSC (%) $\uparrow$ ASD $\downarrow$ DSC (%) $\uparrow$ ASD $\downarrow$ DSC (%) $\uparrow$ ASD $\downarrow$
U-Net (Ronneberger, Fischer, and Brox 2015) 54.82 7.07 77.99 1.16 61.19 8.10 79.49 1.89 68.37 4.56
AmpSwap (Yang and Soatto 2020) 63.56 4.36 78.16 1.13 66.13 4.45 81.84 1.45 72.41 2.85
AmpMix (Xu et al. 2021a) 65.08 4.32 77.06 1.09 68.14 4.29 79.50 1.67 72.44 2.84
FMAug (Li et al. 2023) 63.79 5.03 78.34 1.04 64.33 7.28 82.17 1.29 72.16 3.66
MixStyle (Zhou et al. 2021) 63.95 4.72 80.36 1.06 63.97 6.04 81.93 1.45 72.55 3.32
TriD (Chen et al. 2023) 66.27 4.34 77.08 1.24 65.42 4.53 79.36 1.57 72.03 2.92
ConStyX (Chen et al. 2025) 62.91 5.99 80.69 1.03 64.47 7.16 80.15 1.85 72.05 4.01
RandConv (Xu et al. 2021b) 63.75 4.70 79.35 1.13 65.61 5.07 81.52 1.46 72.56 3.09
CIDA (Ouyang et al. 2022) 53.16 11.34 78.55 5.18 51.27 14.61 68.09 6.62 62.76 8.44
PRandConv (Choi et al. 2023) 44.60 9.84 75.23 1.48 53.34 8.68 76.61 2.57 62.44 5.64
StyCona (ours) 67.80 3.60 78.43 1.08 65.71 4.37 80.68 1.49 73.16 2.63
Figure 3: Qualitative results on the cardiac MRI segmentation and the OC/OD fundus image segmentation tasks.
Results on Cardiac MRI
As reported in Table 1, StyCona achieves consistent performance improvements over the compared methods in the cardiac MRI segmentation task. In general, almost all the compared methods surpass the baseline model, indicating their effectiveness in addressing domain shifts. However, it is worth noting that some augmentation methods, e.g., AmpMix (Xu et al. 2021a) and TriD (Chen et al. 2023), yield segmentation results inferior to those of the baseline model in the bSSFP $\rightarrow$ LGE scenario. This may stem from the fact that the blended amplitude spectrum in AmpMix and the mixed feature statistics in TriD generate style-augmented images that are insufficient for mitigating content shifts, resulting in the models overfitting spurious correlations between the augmented images and their corresponding segmentation labels. In contrast, StyCona demonstrates a clear advantage in handling both style and content shifts by style content augmentation. These results validate the core idea of StyCona and highlight its effectiveness in domain-generalizable medical image segmentation.
Results on OC/OD Fundus Image
In this experiment, we further evaluate StyCona in a single-source, multi-target domain setting. As shown in Table 2, all methods generally achieve comparable average performance across the four target domains. However, the two random convolution-based approaches (i.e., CIDA (Ouyang et al. 2022) and PRandConv (Choi et al. 2023)) yield unsatisfactory results. This may be attributed to that the random convolution operations produce augmented images with distorted content that is inconsistent with the corresponding segmentation labels, thereby misleading the segmentation models during training. In contrast, StyCona demonstrates more robust performance across the four target domains, achieving the highest average DSC and competitive ASD. These results further suggest the effectiveness of StyCona in domain-generalizable medical image segmentation with multiple target domains.
Qualitative results.
As shown in Fig. 3, StyCona produces segmentation results with more precise object delineation. The improvement can be attributed to StyCona’s style and content augmentations, which enable the augmented samples to cover a wide range of unseen domains, thereby allowing the trained model to generate accurate segmentation maps for unseen data. In comparison, methods such as AM (Xu et al. 2021a), MixStyle (Zhou et al. 2021), and PRandConv (Choi et al. 2023) yield relatively unsatisfactory results. AM and MixStyle rely solely on style augmentation by blending amplitude spectra and adjusting the mean and standard deviation of samples, respectively. Meanwhile, the effectiveness of PRandConv is limited by its progressively applied random convolutions, which generate content-distorted images and fail to adequately alleviate variations in anatomical structures. These qualitative results, consistent with the quantitative performance, further suggest the effectiveness of our approach.
Table 3: Ablation study of the proposed method on the cross-sequence setting. The best results are highlighted in bold.
Method StyCona bSSFP $\rightarrow$ LGE LGE $\rightarrow$ bSSFP
  Style augmentation   Content augmentation DSC $\uparrow$ ASD $\downarrow$ DSC $\uparrow$ ASD $\downarrow$
Baseline 65.97 6.55 76.91 3.41
Baseline + Style $\surd$ 68.98 3.20 80.45 2.40
Baseline + Content $\surd$ 72.31 5.16 81.17 2.43
Baseline + Style + Content $\surd$ $\surd$ 73.39 4.33 81.59 2.24
Figure 4: Illustration of a) StyCona-augmented images with different numbers of perturbed content maps ($t=8$, $t=16$, and $t=32$) for two randomly selected original images (The orange dashed circles highlight some content-perturbed regions) and b) Segmentation results for $t=8$, $t=16$, and $t=32$ on cardiac MRI segmentation.
Ablation Study
Effectiveness of each component
In Table 3, we present ablation experiments to analyze the contributions of each component of StyCona. These experiments were conducted under the cardiac MRI segmentation task. The results reveal a consistent trend: segmentation performance improves as the proposed style and content augmentation components are progressively integrated into our method. Specifically, the baseline model (i.e., U-Net) performs relatively unsatisfactorily on the target domain, as it tends to learn domain-specific decision rules based on source domain data. Then, the segmentation performance increases by a large gap when the style augmentation operation is introduced, effectively mitigating style shifts, which are an essential part of domain shifts in these scenarios. In contrast, the improvement is more significant, with an increase of over 7% in DSC, when only the content augmentation component is introduced to address the content shifts, another key aspect of domain shifts. Finally, by incorporating both the style and content augmentation modules to address global variations in image appearance and local differences in anatomical structures, our final method achieves the best performance with DSC of 73.39% and 81.59% on the bSSFP $\rightarrow$ LGE and LGE $\rightarrow$ bSSFP scenarios, respectively. This result demonstrates the effectiveness of each component and validates our assumption that style and content augmentations can alleviate the two types of domain shifts.
Impact of No. Perturbed Content Maps
According to Eq. (1)(2)(3), perturbing the style codes alters only an image’s style without affecting its target shape corresponding to segmentation labels. On the other hand, perturbing $t$ randomly selected content maps results in changes in local anatomical structures (e.g., background tissues or target object-related boundaries), which is the goal of our method to enhance the robustness of segmentation models to anatomical structure variations. Qualitatively, Fig. 4(a) shows the StyCona-augmented samples with varying numbers of perturbed content maps ($t=8,16,32$). Compared with the original images, the augmented images with $t=8,16$ exhibit variations in overall image appearance and regional anatomical structures while remaining consistent with the ground truth segmentation labels. In contrast, at $t=32$, excessive perturbation is introduced into some target boundaries, which may alter their ground truth labels. Quantitatively, among all choices, $t=16$ yields the highest DSC in both the bSSFP $\rightarrow$ LGE and LGE $\rightarrow$ bSSFP scenarios. Based on these results, we set $t=16$ in StyCona to balance perturbation strength with the invariance of segmentation labels, ensuring label consistency after content augmentation.
Conclusion
We propose StyCona, a novel data augmentation algorithm for domain generalizable medical image segmentation. Our findings reveal that an image’s singular values function as style codes, governing its global image properties, while its rank-one matrices serve as content maps, determining its anatomical structures. Based on this decomposition, we categorize domain shifts into style and content shifts w.r.t. deviations in style codes and content maps, respectively. To address these shifts, StyCona perturbs style codes for style augmentation and blends content maps for content augmentation. Extensive experiments on cardiac MRI segmentation with a single target domain and OC/OD fundus image segmentation with multiple target domains demonstrate that StyCona is a promising data augmentation technique for domain-generalizable medical image segmentation.

### v3

@@
-Style Content Decomposition-based Data Augmentation for Domain Generalizable Medical Image Segmentation
+
 Abstract
 Due to domain shifts across diverse medical imaging modalities, learned segmentation models often suffer significant performance degradation during deployment.
-These domain shifts, typically caused by variations in imaging systems, generally comprise two principal components: 1) ”style” shifts, referring to global disparities in image properties such as illumination, contrast, and color; and 2) ”content” shifts, which involve local discrepancies in anatomical structures.
-To address domain shifts in medical image segmentation, a core challenge arises: how can we decouple the factors within images that determine their ”style” and ”content” components?
-To this end, we first propose a linear style-content decomposition method that factorizes an image into style codes and content maps, explicitly modeling the ”style” and ”content” components.
-Building on this, we introduce a Style-Content decomposition-based data augmentation algorithm (StyCona), which leverages this decomposition strategy to guide augmentation of both the global style and local content of source-domain images, enabling the training of a well-generalized model for domain-generalizable medical image segmentation.
+We posit that these domain shifts can generally be categorized into two main components: 1) ”style” shifts, referring to global disparities in image properties such as illumination, contrast, and color; and 2) ”content” shifts, which involve local discrepancies in anatomical structures.
+To address the domain shifts in medical image segmentation, we first factorize an image into style codes and content maps, explicitly modeling the ”style” and ”content” components.
+Building on this, we introduce a Style-Content decomposition-based data augmentation algorithm (StyCona), which performs augmentation on both the global style and local content of source-domain images, enabling the training of a well-generalized model for domain generalizable medical image segmentation.
 StyCona is a simple yet effective plug-and-play module that substantially improves model generalization without requiring additional training parameters or modifications to segmentation model architectures.
 Experiments on cardiac magnetic resonance imaging and fundus photography segmentation tasks, with single and multiple target domains respectively, demonstrate the effectiveness of StyCona and its superiority over state-of-the-art domain generalization methods.
-The code will be released at https://github.com/Senyh/StyCona.
-Introduction
+The code is available at https://github.com/Senyh/StyCona.
+keywords:
+Medical Image Segmentation, Domain Generalization, Style Code, Content Map
+1 Introduction
 Medical image segmentation is critical for computer-aided diagnosis.
-Driven by large and diverse annotated datasets, deep learning-based segmentation models have achieved remarkable progress (Ronneberger, Fischer, and Brox 2015; Milletari, Navab, and Ahmadi 2016).
-However, data distribution mismatches (a.k.a., domain shifts), caused by diverse imaging protocols, equipment vendors, or patient populations, etc., between source (training) and target (testing) domains hinder the generalizability of trained models for clinical deployment (Castro, Walker, and Glocker 2020; Zhou et al.
-2022; Guan and Liu 2021).
-Figure 1: Illustration of style codes and content maps for bSSFP and LGE MRI sequences: a) original images showing disparities in both style (global image appearance) and content (local anatomical structures), b) image histograms reflecting domain shifts in terms of pixel intensities, c) style-swapped images generated by swapping the two original images’ style codes (top: bSSFP image with LGE style; bottom: LGE image with bSSFP style), d) content maps of the original images, and e) t-SNE (Van der Maaten and Hinton 2008) visualization of domain shifts (top: style shift; bottom: content shift).
-Domain generalization (DG) has emerged as a promising strategy to mitigate domain shifts by training models solely on source-domain data and generalizing them to unseen domains.
-Current domain generalization (DG) methods encourage segmentation models to learn domain-invariant representations through representation learning (Chen et al.
-2019, 2020; Pei et al.
-2021; Peng et al.
-2022; Gao et al.
-2022, 2023), data augmentation (a.k.a., domain randomization) (Ouyang et al.
-2022; Zhou et al.
-2021; Chen et al.
-2023, 2025), as well as the combination of both schemes (Gu et al.
-2023; Yang and Soatto 2020; Xu et al.
-2021a; Li et al.
-2023).
-Specifically, the first line of work focuses on explicitly learning domain-invariant information via deterministic (Chen et al.
-2019, 2020; Pei et al.
-2021) or statistical (Gao et al.
-2022, 2023) modeling.
-While these methods yield intuitive decoupled features on source-domain data, domain shifts pose a dilemma: the disentanglement models, constrained by domain-specific training, often fail to generalize their disentanglement capabilities across different domains.
-In contrast, data augmentation-based DG methods can circumvent this issue by expanding the source-domain data distribution and thus implicitly encouraging models to capture domain-invariant information.
-These methods leverage techniques such as random convolution (Xu et al.
-2021b; Ouyang et al.
-2022; Choi et al.
-2023), and feature statistics editing (Zhou et al.
-2021; Chen et al.
-2023, 2025), etc.
-The third line of research employs learning-based (Gu et al.
-2023) or rule-based (Yang and Soatto 2020; Xu et al.
-2021a; Li et al.
-2023) disentanglement to decompose images into domain-invariant contents and domain-specific styles, followed by style augmentation based on the disentangled components.
-However, the former depends on additional encoders and decoders, as well as access to multiple domain images and their corresponding labels.
-These requirements are computationally intensive and may be infeasible in practice, e.g., single-source domain generalization or domain-label unavailable scenarios.
-In contrast, the latter assumes invariance in the Fourier phase spectrum, which limits its effectiveness for data augmentation to alleviate domain shifts.
-Moreover, the specific components contributing to domain shifts remain underexplored.
-This raises a question: Can a deeper understanding of the underlying nature of domain shifts guide the design of more effective data augmentation methods for domain generalizable medical image segmentation?
-To answer this question, we begin with a visual investigation of domain shifts in medical images.
-As shown in Fig.1(a-b), domain shifts generally manifest as variations in pixel intensity, comprising two major components: 1) style shift, referring to global disparities in image characteristics such as brightness, contrast, and color; and 2) content shift, referring to local anatomical discrepancies caused by differences in tissue visualization across imaging systems.
-Inspired by these observations, rather than learning-based (Gu et al.
-2023) or rule-based (Yang and Soatto 2020; Xu et al.
-2021a; Li et al.
-2023) disentanglement, we explicitly model the ”style” and ”content” components of an image by linearly decomposing it into a style code and content maps.
-We reveal that 1) the style code captures global image characteristics within a given anatomical structure and 2) the content maps describe the image’s anatomical structure.
-To provide a clearer illustration, we visualize style-swapped images (by exchanging two images’ style codes) and content maps in Fig.
-1(c-d); it can be observed that the content maps delineate the anatomical structures of the original bSSFP and LGE images, and the style-swapped images transfer appearance between each other while preserving the original anatomical structures.
+Driven by large and diverse annotated datasets, deep learning-based segmentation models have achieved remarkable progress (ronneberger2015u; milletari2016v).
+However, data distribution mismatches (a.k.a., domain shifts), caused by diverse imaging protocols, equipment vendors, or patient populations, etc., between source (training) and target (testing) domains hinder the generalizability of trained models for clinical deployment (castro2020causality; zhou2022domain; guan2021domain).
+Figure 1: Illustration of style codes and content maps for bSSFP and LGE MRI sequences: a) original images showing disparities in both style (global image appearance) and content (local anatomical structures), b) image histograms reflecting domain shifts in terms of pixel intensities, c) style-swapped images generated by swapping the two original images’ style codes (top: bSSFP image with LGE style; bottom: LGE image with bSSFP style), d) content maps of the original images, and e) t-SNE (van2008visualizing) visualization of domain shifts (top: style shift; bottom: content shift).
+Domain generalization (DG) aims to mitigate domain shifts by training models on source-domain data and generalizing them to unseen target domains.
+Existing DG studies generally promote the learning of domain-invariant representations through two major paradigms: representation learning (chen2019synergistic; chen2020unsupervised; pei2021disentangle; peng2022unsupervised; gao2022joint; gao2023bayeseg) and data augmentation (a.k.a., domain randomization) (ouyang2022causality; zhou2021mixstyle; chen2023treasure; chen2025constyx; gu2023cddsa; yang2020fda; xu2021fourier; li2023frequency).
+The former explicitly explores domain-invariant information via deterministic (chen2019synergistic; chen2020unsupervised; pei2021disentangle) or statistical (gao2022joint; gao2023bayeseg) modeling.
+Although such methods yield intuitive decoupled features on source-domain data, their reliance on domain-specific training often hinders the generalization of disentanglement capabilities across different domains.
+In contrast, data augmentation-based DG methods alleviate this limitation by expanding the source-domain data distribution and thus implicitly encouraging models to excavate domain-invariant features.
+As a result, this paradigm has become the mainstream approach in the field.
+Specifically, these methods typically employ techniques such as Fourier transformation (yang2020fda; xu2021fourier; li2023frequency), random convolution (xu2021robust; ouyang2022causality; choi2023progressive), and feature statistics editing (zhou2021mixstyle; chen2023treasure; chen2025constyx), etc to achieve style augmentation on source-domain data.
+However, the underlying components involved in medical images remain underexplored.
+Modeling these components can provide valuable insights into the nature of domain shifts and guide the design of more effective data augmentation strategies.
+To bridge this gap, we introduce a style–content decomposition strategy that decomposes an image into style codes and content maps, revealing that 1) the style code captures global image characteristics within a given anatomical structure and 2) the content maps describe the image’s anatomical structure.
+Fig.
+1(c-d) visualizes style-swapped images (obtained by exchanging the style codes between two images) and content maps.
+It can be observed that the content maps delineate the anatomical structures of the original bSSFP and LGE images, while the style-swapped images exchange appearance between domains but preserve the original anatomical structures.
 Furthermore, the shift observed in the style codes is more significant than that in the content maps [Fig.
-1(e)], as global variations in image characteristics are generally more substantial than local differences in anatomical structures.
-These results align with the aforementioned intuition.
-Based on this, we categorize domain shifts into two components: 1) style shifts (global image property variations) as indicated by deviations in the style codes, and 2) content shifts (local anatomical structure discrepancies) as reflected in differences between the content maps.
-Given that explicitly measuring and reducing these two types of shifts is infeasible due to the unavailability of target domain data during training, we instead aim to perturb the ”style” and ”content” components of source domain images for generating images from diverse domains (simulating patients undergoing different imaging systems) to train a domain generalizable medical image segmentation model.
+1(e)], as global variations in image characteristics are generally greater than local differences in anatomical structures.
+Based on these observations, we categorize domain shifts in medical images into two components: 1) style shifts (global image property variations) as indicated by deviations in the style codes, and 2) content shifts (local anatomical structure discrepancies) as reflected in differences between the content maps.
+Since quantitatively measuring and reducing these two types of shifts is infeasible without access to target domain data during training, we instead perturb the ”style” and ”content” components of source domain images to generate augmented images from diverse domains (simulating patients undergoing different imaging systems), enabling the training of a well-generalized medical image segmentation model.
 To this end, we propose a style content decomposition-based data augmentation algorithm (StyCona) to advance domain generalizable medical image segmentation.
-Specifically, StyCona performs perturbations on the ”style” (global image characteristics) and ”content” (local anatomical structures) by 1) blending style codes and 2) mixing content maps.
+Specifically, StyCona performs perturbations on the ”style” (global image characteristics) and ”content” (local anatomical structures) of an image by 1) blending style codes and 2) mixing content maps, respectively.
+StyCona generates augmented images with diverse styles and contents while preserving their semantic information, enabling the simulation of images from a wide range of domains for training a well-generalized segmentation model.
 We evaluate StyCona on cross-domain cardiac magnetic resonance imaging (MRI) segmentation and optic cup (OC)/optic disk (OD) fundus photography segmentation tasks.
-Experimental results demonstrate that StyCona is a promising solution for domain generalizable medical image segmentation and achieves state-of-the-art performance.
+Experimental results demonstrate that StyCona is a promising solution for domain generalizable medical image segmentation.
 Our main contributions can be summarized as follows:
   • $\bullet$
-    To the best of our knowledge, this is the first work that decomposes domain shifts into style shifts and content shifts through linear decomposition of images in domain generalizable medical image segmentation tasks.
+    We propose a style–content decomposition strategy to model the underlying components of medical images and provide a deep insight into domain shifts.
   • $\bullet$
     We propose StyCona, a novel data augmentation algorithm for domain generalizable medical image segmentation.
-Our method can be easily integrated into off-the-shelf medical image segmentation backbones and effectively mitigates the two types of domain shifts.
+StyCona can be easily integrated into off-the-shelf medical image segmentation backbones and effectively mitigates domain shifts.
   • $\bullet$
     StyCona leads to state-of-the-art performance across a wide range of tasks and provides a compelling alternative to Fourier transformation-based, random convolution-based, and feature statistics editing-based domain generalization methods.
-Related work
-Domain Generalization
-Domain generalization (DG) aims to train a model on a source domain and enable it to generalize to unseen target domains.
-Numerous DG methods have been proposed for both natural and medical image scenarios (Zhou et al.
-2022; Castro, Walker, and Glocker 2020).
-Specifically, these methods enforce the learning of domain-invariant representations through various approaches, including: deterministic or statistical modeling of domain-invariant and domain-specific features (Gu et al.
-2023; Gao et al.
-2022, 2023); frequency transformation-based mix-up, which combines Fourier amplitude spectra from different images for style augmentation (Yang and Soatto 2020; Xu et al.
-2021a; Li et al.
-2023; Zhao et al.
-2024); random convolution-based data augmentation, which perturbs pixel intensities via convolution neural networks with random parameters (Xu et al.
-2021b; Ouyang et al.
-2022; Choi et al.
-2023); as well as feature statistics editing-based style transfer, which perturbs image- or feature-level statistics (e.g., mean and standard deviation) to achieve style or content augmentation (Zhou et al.
-2021; Chen et al.
-2023, 2025).
-These techniques are generally applicable to both image recognition and semantic segmentation tasks, with the latter requiring additional consideration for transformation equivalence.
-However, the fundamental nature of domain shifts, crucial for guiding the development of more effective DG approaches, remains underexplored.
-In this paper, we propose a novel framework that first decomposes domain shifts into style and content shifts through linear modeling of images and then designs a novel data augmentation algorithm to mitigate both types of domain shifts.
-Content-Style Decomposition
-Content-style decomposition disentangles features into domain-invariant content and domain-specific style components for addressing the domain shift issue.
-In the context of semantic segmentation, this technique has been widely employed in the design of both unsupervised domain adaptation (Chen et al.
-2019, 2020; Pei et al.
-2021; Peng et al.
-2022; Yang and Soatto 2020) and domain generalization (Gu et al.
-2023; Gao et al.
-2022, 2023; Xu et al.
-2021a; Li et al.
-2023; Zhao et al.
-2024) frameworks.
-Research in this area generally falls into two categories: 1) learning-based (Chen et al.
-2019, 2020; Pei et al.
-2021; Peng et al.
-2022; Gu et al.
-2023; Gao et al.
-2022, 2023) and 2) rule-based approaches (Yang and Soatto 2020; Xu et al.
-2021a; Li et al.
-2023; Zhao et al.
-2024).
-Learning-based methods rely on additional deep neural network-based encoders and decoders, with specific constraints (e.g., contrastive learning on decoupled components (Gu et al.
-2023)), to achieve feature disentanglement either deterministically (Chen et al.
-2019, 2020; Pei et al.
-2021; Peng et al.
-2022; Gu et al.
-2023) or statistically (Gao et al.
-2022, 2023).
-Rule-based methods adopt the Fourier transformation to perform direct decomposition.
-However, they impose the constraint that phase spectra should be fixed to ensure the invariance of the original contents during reconstruction.
-This constraint limits their flexibility and effectiveness for data augmentation.
-In contrast, we propose a simple yet effective content-style decomposition method that linearly factorizes an image into a style code and content maps, enabling style-content augmentation through perturbations on the decoupled components.
 Figure 2: Schematic diagram of the proposed style content decomposition-based data augmentation algorithm (StyCona).
-StyCona includes three steps: 1) singular value decomposition, 2) style augmentation, and 3) content augmentation.
+StyCona includes three steps: 1) style content decomposition, 2) style augmentation, and 3) content augmentation.
 $x_{i}$ and $x_{j}$ represent original images, and $\bar{\tilde{x}}$ denotes a StyCona-augmented sample.
-Methodology
+2 Methodology
 For a domain generalizable medical image segmentation task, the training set can be denoted as $\mathcal{D}=\{\mathcal{S}_{k}|\mathcal{S}_{k}=\{(x_{i(k)},y_{i(k)})_{i=1}^{N_{k}}\}_{k=1}^{K}\}$, where $x_{i(k)}$ represents the $i^{th}$ image in the $k^{th}$} domain and $y_{i(k)}$ is its corresponding ground truth segmentation label.
@@
 Note that we omit the domain index for brevity.
-Overview
-We assume that domain shifts in medical images resulting from variations in image acquisition processes generally comprise two principal components: 1) ”style” shifts, referring to global variations in image properties such as illumination, contrast, and color; and 2) ”content” shifts, which involve local changes in image characteristics and are usually associated with differences in the visibility of specific tissues.
-Under this assumption, a key problem arises for addressing the domain shifts: how can we decouple the factors within images that determine the ”style” and ”content” components?
-Although several approaches have been proposed to address this problem, they either rely on auxiliary disentanglement networks (Gu et al.
-2023; Gao et al.
-2022, 2023) or impose constraints on the perturbation of the decoupled components (Xu et al.
-2021a).
-Motivated by the effectiveness of linear decomposition in image compression (Andrews and Patterson 2003; Ranade, Mahabalarao, and Kale 2007) and restoration (Zhu et al.
-2014; Guo et al.
-2015; Gao and Zhuang 2020), we propose to linearly decompose an image into its style code and content maps, from which we validate both aforementioned types of domain shifts, as shown in Fig.
-1.
-Based on the linear decomposition, we introduce a style content decomposition-based data augmentation algorithm, termed StyCona, for domain generalizable medical image segmentation.
-The training process for leveraging StyCona to train a domain-generalizable medical image segmentation model is summarized in Algorithm 1.
-Style Content Decomposition-based Data Augmentation (StyCona)
-The core idea behind StyCona is that the style and content shifts can be mitigated through style and content augmentations, respectively.
-To this end, the StyCona algorithm mainly includes two components: 1) image decomposition and 2) style content augmentation.
-Specifically, the image decomposition is achieved using singular value decomposition (SVD) (Klema and Laub 1980) to factorize an image into its singular values (style codes) and rank-one matrices (content maps) The style content augmentation is then performed by perturbing style codes and content maps, respectively.
+2.1 Style Content Decomposition-based Data Augmentation (StyCona)
+Domain shifts in medical images can be characterized two principal components: 1) ”style” shifts, referring to global variations in image properties such as illumination, contrast, and color; and 2) ”content” shifts, which involve local changes in image characteristics and are usually associated with differences in the visibility of specific tissues.
+Based on this assumption, we propose StyCona to mitigate both style and content shifts through corresponding style and content augmentations.
 As illustrated in Fig.
-2, we realize the StyCona algorithm with three steps: 1) singular value decomposition, 2) style augmentation, and 3) content augmentation.
-Singular Value Decomposition
-For an image $x$, its SVD is formulated as:
-$x=\sum_{k}\underbrace{\sigma^{k}}_{style\,code}\,\,\underbrace{\boldsymbol{u}^{k}{\boldsymbol{v}^{k}}^{\mathsf{T}}}_{content\,map}$ (1)
-where each singular value $\sigma^{k}$ acts as a style code and the rank-one matrix $\boldsymbol{u}^{k}{\boldsymbol{v}^{k}}^{\mathsf{T}}$ serves as a content map for the subsequent style and content augmentation.
-Algorithm 1 StyCona for domain generalizable medical image segmentation
-Input: $\mathcal{D}=\{(x_{i},y_{i})_{i=1}^{N}\}$
-Parameter: $\theta$ for the segmentation model $f(\cdot;\theta)$
-Output: Trained model $f(\cdot;\theta)$
-1:for each iteration do
-2:  # 1) Extract style codes and content maps from current batch original image $x_{i}$ and auxiliary image $x_{j}$ ($i\neq j$) by Eq. (1)
-3:  $\boldsymbol{\sigma}_{i}\leftarrow x_{i}\,\,and\,\,\{\boldsymbol{u}^{r}_{i}\times\boldsymbol{v}^{r}_{i}\}\leftarrow x_{i}$
-4:  $\boldsymbol{\sigma}_{j}\leftarrow x_{j}\,\,and\,\,\{\boldsymbol{u}^{r}_{j}\times\boldsymbol{v}^{r}_{j}\}\leftarrow x_{i}$
-5:  # 2) Style Augmentation by Eq. (2)
-6:  $\tilde{\sigma}_{i}^{k}=\alpha\times\sigma_{i}^{k}+(1-\alpha)\times\sigma_{j}^{k}$
-7:  # 3) Content Augmentation by Eq. (3)
-8:  $\bar{\boldsymbol{u}}_{i}^{k}=\beta\times\boldsymbol{u}_{i}^{k}+(1-\beta)\times\boldsymbol{u}_{j}^{k}\quad or\quad\bar{\boldsymbol{v}}_{i}^{k}=\beta\times\boldsymbol{v}_{i}^{k}+(1-\beta)\times\boldsymbol{v}_{j}^{k}$
-9:  # 4) Obtain StyCona images via inverse singular value decomposition
-10:  $\bar{\tilde{x}}_{i}=\sum^{r_{i}}_{k=1}\tilde{\boldsymbol{\sigma}}_{i}^{k}\bar{\boldsymbol{u}}_{i}^{k}{\bar{\boldsymbol{v}}_{i}^{k}}{}^{\mathsf{T}}$
-11:  # 5) Forward propagation to obtain the predictive label for the StyCona image $\bar{\tilde{x}}_{i}$
-12:  $\hat{\bar{\tilde{y}}}_{i}\leftarrow f(\bar{\tilde{x}}_{i},\theta)$
-13:  # 6) Loss supervision by Eq. (4)
-14:  Back-Propagate $\mathcal{L}$
-15:  Update $\theta$
-16:  return $f(\cdot;\theta)$
-17:end for
-Style Augmentation
+2, StyCona mainly includes three steps: 1) style content decomposition, 2) style augmentations, and 3) content augmentation.
+Specifically, the style-content decomposition is performed using singular value decomposition (SVD) (klema1980singular), which factorizes an image into its singular values (style codes) and rank-one matrices (content maps).
+Then, the style and content augmentations are achieved by perturbing the style codes and content maps, respectively.
+2.1.1 Style Content Decomposition
+Low-rank matrices of an image correspond to its principal components abdi2010principal; candes2011robust.
+Based on this idea, prior work has employed rank-one matrices as the principal parts of an image for restoration tasks gao2020rank.
+We leverage SVD to show that an image’s singular values encode its style codes, while the associated rank-one matrices represent its content maps:
+$x=\sum^{R}_{r}\underbrace{\sigma^{r}}_{style\,code}\,\,\underbrace{\boldsymbol{u}^{r}{\boldsymbol{v}^{r}}^{\mathsf{T}}}_{content\,map}$ (1)
+where $R$ denotes the rank of the image, corresponding to the number of non-zero singular values.
+SVD provides the optimal rank-one decomposition in the least-squares sense.
+It guarantees that the set $\boldsymbol{\Phi}=\{\boldsymbol{u}^{1}{\boldsymbol{v}^{1}}^{\mathsf{T}},\boldsymbol{u}^{2}{\boldsymbol{v}^{2}}^{\mathsf{T}},...,\boldsymbol{u}^{R}{\boldsymbol{v}^{R}}^{\mathsf{T}}\}$ forms a complete basis for the column and row spaces of $x$.
+  • Content Map.
+Each rank-one matrix in $\boldsymbol{\Phi}$ functions as a basis pattern of the image, and any image structure can be represented as a linear combination of these patterns.
+Thus, $\boldsymbol{\Phi}$ completely defines the anatomical structures of the image.
+  • Style Code.
+Each scalar $\sigma^{r}$ acts as a global multiplier for the corresponding basis pattern $(\boldsymbol{u}^{r}{\boldsymbol{v}^{r}}^{\mathsf{T}})$.
+It modulates the magnitude of pixel intensity in the image, scaling the values of the basis pattern proportionally.
+As illustrated in Fig.
+1, swapping the style codes of the bSSFP and LGE images results in style transfer between them, while summing all content maps of one of the images reconstructs its complete anatomical structure.
+2.1.2 Style Augmentation
 Style augmentation aims to alleviate style shifts (global variations in image properties).
 As depicted in Fig.
-2, we perform style augmentation on an image $x_{i}$ by blending its style codes (i.e., singular values) with those of an auxiliary image $x_{j}$ in an element-wise manner:
-$\tilde{\sigma}_{i}^{k}=\alpha\times\sigma_{i}^{k}+(1-\alpha)\times\sigma_{j}^{k}$ (2)
-where the weight $\alpha\sim U(0,1)$ controls the mixing strength, $\sigma_{i}^{k}$ represents the $k_{th}$ singular value of image $x_{i}$, $i/j$ indicates the sample index and $i\neq j$.
-After that, a style-augmented image $\tilde{x}_{i}$ can be generated using the perturbed singular values and the original singular vectors by Eq.
-(1).
-Content Augmentation
-We devise image content augmentation to mitigate content shifts (local disparities in anatomical structure).
-This is achieved by mixing a set of $t$ randomly selected content maps (i.e., rank-one matrices) from an image $x_{i}$ with another set of $t$ randomly selected content maps from an auxiliary image $x_{j}$ [Fig.
+2, we perform style augmentation on an image $x_{i}$ by blending its style codes with those of an auxiliary image $x_{j}$ in an element-wise manner:
+$\tilde{\sigma}_{i}^{r}=\alpha\times\sigma_{i}^{r}+(1-\alpha)\times\sigma_{j}^{r}$ (2)
+where the weight $\alpha\sim U(0,1)$ controls the mix strength, $\sigma_{i}^{r}$ represents the $r_{th}$ singular value of image $x_{i}$, and $i/j$ indicates the sample index ($i\neq j$).
+2.1.3 Content Augmentation
+We devise content augmentation to mitigate content shifts (local disparities in anatomical structure).
+This is achieved by mixing a set of $t$ randomly selected content maps from an image $x_{i}$ with another set of $t$ randomly selected content maps from an auxiliary image $x_{j}$ [Fig.
 2].
 The perturbation to each content map is formulated as:
-$\bar{\boldsymbol{u}}_{i}^{k}=\beta\times\boldsymbol{u}_{i}^{k}+(1-\beta)\times\boldsymbol{u}_{j}^{k}\quad or\quad\bar{\boldsymbol{v}}_{i}^{k}=\beta\times\boldsymbol{v}_{i}^{k}+(1-\beta)\times\boldsymbol{v}_{j}^{k}$ (3)
-where the weight $\beta\sim U(0,1)$ and $\boldsymbol{u}_{i/j}^{k}$/$\boldsymbol{v}_{i/j}^{k}$ represents the left/right singular vector of image $x_{i/j}$.
-Using the perturbed content maps and the original style codes, we can obtain a content-augmented image $\bar{x}_{i}$ via Eq.
-(1).
-In a nutshell, by leveraging Eq.
-(1)(2)(3), StyCona conducts style and content augmentation and produces an augmented image $\bar{\tilde{x}}_{i}=\sum^{r_{i}}_{k=1}\tilde{\boldsymbol{\sigma}}_{i}^{k}\bar{\boldsymbol{u}}_{i}^{k}{\bar{\boldsymbol{v}}_{i}^{k}}{}^{\mathsf{T}}$.
-Loss Supervision
-Afterwards, we formulate the segmentation loss $\mathcal{L}$ for training a domain generalizable model $f(\cdot;\theta)$ based on StyCona images:
+$\bar{\boldsymbol{u}}_{i}^{r}=\beta\times\boldsymbol{u}_{i}^{r}+(1-\beta)\times\boldsymbol{u}_{j}^{r}\quad or\quad\bar{\boldsymbol{v}}_{i}^{r}=\beta\times\boldsymbol{v}_{i}^{r}+(1-\beta)\times\boldsymbol{v}_{j}^{r}$ (3)
+where the weight $\beta\sim U(0,1)$ and $\boldsymbol{u}_{i/j}^{r}$$\boldsymbol{v}_{i/j}^{r}$ represents the left/right singular vector of image $x_{i/j}$.
+In a nutshell, StyCona performs style and content augmentation by Eq.
+(1)(2)(3), and generates an augmented image as $\bar{\tilde{x}}_{i}=\sum^{R}_{r=1}\tilde{\boldsymbol{\sigma}}_{i}^{r}\bar{\boldsymbol{u}}_{i}^{r}{\bar{\boldsymbol{v}}_{i}^{r}}{}^{\mathsf{T}}$.
+2.1.4 Loss Supervision
+Afterwards, we formulate the segmentation loss $\mathcal{L}$ for training a domain generalizable model $f(\cdot;\theta)$ based on StyCona augmented images:
 $\mathcal{L}=\frac{1}{N}\sum^{N}_{i}\mathcal{L}_{seg}(f(\bar{\tilde{x}}_{i},\theta),y_{i})$ (4)
 where $\mathcal{L}_{seg}$ denotes a segmentation criterion.
-Experiments and Results
+3 Experiments and Results
 Cross-Domain Cardiac Magnetic Resonance Imaging (MRI) Segmentation
-The MS-CMR Dataset (Single Source and Single Target Domain) dataset (Zhuang 2018) contains 45 subjects, each with bSSFP and LGE MRI sequences, along with ground truth annotations for the right ventricle (RV), left ventricle (LV), and myocardium (MYO).
-All images are normalized to the range [0,1][0,1][ 0 , 1 ] and resampled to a uniform resolution of $\times 1.0$ 1.0 mm.
+The MS-CMR Dataset (Single Source and Single Target Domain) dataset (zhuang2018multivariate) contains 45 subjects, each with bSSFP and LGE MRI sequences, along with ground truth annotations for the right ventricle (RV), left ventricle (LV), and myocardium (MYO).
+All images are normalized to the range $[0,1]$ and resampled to a uniform resolution of $1.0\times 1.0$ mm.
 This experimental setting evaluates a segmentation model’s ability to generalize across MRI sequences for cardiac structure segmentation, specifically from LGE to bSSFP and vice versa.
-For each direction, one sequence (e.g., bSSFP) is treated as the source domain and is divided into training and validation sets with a 8:28:28 : 2 ratio, while another sequence (e.g., LGE) serves as the target domain and is used solely for testing.
+For each direction, one sequence (e.g., bSSFP) is treated as the source domain and is divided into training and validation sets with a $8:2$ ratio, while another sequence (e.g., LGE) serves as the target domain and is used solely for testing.
 Cross-Domain Optic Cup (OC) and Optic Disc (OD) Fundus Image Segmentation
-The Fundus Image Benchmark (Single Source and Multiple Target Domains(Chen et al.
-2023) is established based on the ORIGA (Zhang et al.
-2010), Drishti-GS (Sivaswamy et al.
-2014), REFUGE (Orlando et al.
-2020), and RIGA (Almazroa et al.
-2018) (BinRushed and Magrabia) datasets.
+The Fundus Image Benchmark (Single Source and Multiple Target Domains(chen2023treasure) is established based on the ORIGA (zhang2010origa), Drishti-GS (sivaswamy2014drishti), REFUGE (orlando2020refuge), and RIGA (almazroa2018retinal) (BinRushed and Magrabia) datasets.
 This setting evaluates the model’s generalization capability for the joint OC and OD segmentation across multiple target domains.
-The ORIGA dataset is used as the source domain with the training and validation split following TriD (Chen et al.
-2023); the other datasets (BinRushed, Drishti-GS, Magrabia, and REFUGE) serve as target domains for cross-domain evaluation.
+The ORIGA dataset is used as the source domain with the training and validation split following TriD (chen2023treasure); the other datasets (BinRushed, Drishti-GS, Magrabia, and REFUGE) act as target domains for cross-domain evaluation.
 Implementation Details.
-We conducted the experiments using PyTorch (Paszke et al.
-2019) on an NVIDIA A40 GPU with 48G GPU memory.
-All compared methods were optimized using an AdamW optimizer (Kingma and Ba 2015) with a fixed learning rate of $1e-4$ during the 100 training epochs for both the cadiac MRI and fundus image segmentation tasks.
-U-Net (Ronneberger, Fischer, and Brox 2015) was employed as the baseline segmentation model.
-The combination of cross-entropy and Dice loss (Milletari, Navab, and Ahmadi 2016) was used as the segmentation criterion.
-All images were resized to $\times 256$ 256 for both training and testing.
-During evaluation, the predicted labels were rescaled to their original resolutions.
-We set the number of perturbed content maps $t=16$ (please refer to the ablation study for further analysis).
-Comparison with State of the Arts
-We extensively compared our StyCona with state-of-the-art DG methods: 1) Fourier transformation-based: amplitude swap (AmpSwap) (Yang and Soatto 2020), amplitude mixup (AmpMix) (Xu et al.
-2021a), and FMAug (Li et al.
-2023)); 2) Feature statistics editing-based: MixStyle (Zhou et al.
-2021), TriD (Chen et al.
-2023), and ConStyX (Chen et al.
-2025); 3) Random convolution-based: RandConv (Xu et al.
-2021b), CIDA (Ouyang et al.
-2022), and PRandConv (Choi et al.
-2023).
+We conducted the experiments using PyTorch (paszke2019pytorch) on an NVIDIA A40 GPU with 48G GPU memory.
+All compared methods were optimized using an AdamW optimizer (kingma2014adam) with a fixed learning rate of 1e41e-4 during the 100 training epochs for both the cadiac MRI and fundus image segmentation tasks.
+U-Net (ronneberger2015u) was employed as the baseline segmentation model.
+The combination of cross-entropy and Dice loss (milletari2016v) was used as the segmentation criterion.
+All images were resized to $256\times 256$ for both training and testing, and the predicted labels were rescaled to their original resolutions for evaluation.
+We set the number of perturbed content maps $t=16$ (please refer to Section 3.2.2 for further analysis).
+3.1 Comparison with State of the Arts
+We compared StyCona with state-of-the-art DG methods: 1) Fourier transformation-based: amplitude swap (AmpSwap) (yang2020fda), amplitude mixup (AmpMix) (xu2021fourier), and FMAug (li2023frequency)); 2) Feature statistics editing-based: MixStyle (zhou2021mixstyle), TriD (chen2023treasure), and ConStyX (chen2025constyx); 3) Random convolution-based: RandConv (xu2021robust), CIDA (ouyang2022causality), and PRandConv (choi2023progressive).
 We re-implemented all the compared methods in a unified experimental setup for fair comparison.
@@
 3], showcasing its superiority in domain generalizable medical image segmentation.
+3.1.1 Results on Cardiac MRI
+As reported in Table 1, StyCona achieves consistent performance improvements over the compared methods in the cardiac MRI segmentation task.
+In general, almost all the compared methods surpass the baseline model, indicating their effectiveness in addressing domain shifts.
+However, it is worth noting that some methods, e.g., AmpMix (xu2021fourier) and TriD (chen2023treasure), yield segmentation results inferior to those of the baseline model in the bSSFP $\rightarrow$ LGE scenario.
+This may result from the fact that the blended amplitude spectrum in AmpMix and the mixed feature statistics in TriD generate style-augmented images that are insufficient for mitigating content shifts, resulting in the models overfitting spurious correlations between the augmented images and their corresponding segmentation labels.
+In contrast, StyCona demonstrates a clear advantage in handling both style and content shifts by style content augmentation.
+These results validate the key idea of StyCona and highlight its effectiveness in domain generalizable medical image segmentation.
+3.1.2 Results on Fundus Image
+We further evaluate StyCona in a single-source, multi-target domain setting.
+As shown in Table 2, all methods generally achieve comparable average performance across the four target domains.
+However, the two random convolution-based approaches (i.e., CIDA (ouyang2022causality) and PRandConv (choi2023progressive)) yield relatively unsatisfactory results.
+This may be attributed to that the random convolution operations produce augmented images with distorted content that is inconsistent with the corresponding segmentation labels, thus misleading the segmentation models during training.
+In contrast, StyCona shows more robust performance across the four target domains, obtaining the highest average DSC and competitive ASD.
+These results further suggest the effectiveness of StyCona in domain generalizable medical image segmentation with multiple target domains.
 Table 1: Comparison with state-of-the-art methods on the cross-domain cardiac MRI segmentation task (single target domain).
@@
 DSC $\uparrow$ ASD $\downarrow$ DSC $\uparrow$ ASD $\downarrow$
-U-Net (Ronneberger, Fischer, and Brox 2015) 65.97 6.55 76.91 3.41
-AmpSwap (Yang and Soatto 2020) 70.93 6.48 78.74 3.26
-AmpMix (Xu et al. 2021a) 63.89 6.44 80.07 2.46
-FMAug (Li et al. 2023) 72.46 13.67 80.54 2.57
-MixStyle (Zhou et al. 2021) 69.26 10.91 80.81 2.69
-TriD (Chen et al. 2023) 65.12 10.25 79.14 4.60
-ConStyX (Chen et al. 2025) 68.63 7.89 81.30 2.85
-RandConv (Xu et al. 2021b) 71.90 5.65 75.17 4.27
-CIDA (Ouyang et al. 2022) 69.72 8.63 72.85 4.61
-PRandConv (Choi et al. 2023) 56.61 8.75 75.57 3.38
+U-Net (ronneberger2015u) 65.97 6.55 76.91 3.41
+AmpSwap (yang2020fda) 70.93 6.48 78.74 3.26
+AmpMix (xu2021fourier) 63.89 6.44 80.07 2.46
+FMAug (li2023frequency) 72.46 13.67 80.54 2.57
+MixStyle (zhou2021mixstyle) 69.26 10.91 80.81 2.69
+TriD (chen2023treasure) 65.12 10.25 79.14 4.60
+ConStyX (chen2025constyx) 68.63 7.89 81.30 2.85
+RandConv (xu2021robust) 71.90 5.65 75.17 4.27
+CIDA (ouyang2022causality) 69.72 8.63 72.85 4.61
+PRandConv (choi2023progressive) 56.61 8.75 75.57 3.38
 StyCona (ours) 73.39 4.33 81.59 2.24
-Table 2: Comparison with state-of-the-art methods on the fundus segmentation tasks with 10% labeled data (Source: ORIGA; Target: REFUGE, Drishti-GS, BinRushed, and Magrabia).
+Table 2: Comparison with state-of-the-art methods on the fundus image segmentation tasks (Source: ORIGA; Target: REFUGE, Drishti-GS, BinRushed, and Magrabia).
 The best and second-best results are highlighted in bold and underline, respectively.
@@
 DSC (%) $\uparrow$ ASD $\downarrow$ DSC (%) $\uparrow$ ASD $\downarrow$ DSC (%) $\uparrow$ ASD $\downarrow$ DSC (%) $\uparrow$ ASD $\downarrow$ DSC (%) $\uparrow$ ASD $\downarrow$
-U-Net (Ronneberger, Fischer, and Brox 2015) 54.82 7.07 77.99 1.16 61.19 8.10 79.49 1.89 68.37 4.56
-AmpSwap (Yang and Soatto 2020) 63.56 4.36 78.16 1.13 66.13 4.45 81.84 1.45 72.41 2.85
-AmpMix (Xu et al. 2021a) 65.08 4.32 77.06 1.09 68.14 4.29 79.50 1.67 72.44 2.84
-FMAug (Li et al. 2023) 63.79 5.03 78.34 1.04 64.33 7.28 82.17 1.29 72.16 3.66
-MixStyle (Zhou et al. 2021) 63.95 4.72 80.36 1.06 63.97 6.04 81.93 1.45 72.55 3.32
-TriD (Chen et al. 2023) 66.27 4.34 77.08 1.24 65.42 4.53 79.36 1.57 72.03 2.92
-ConStyX (Chen et al. 2025) 62.91 5.99 80.69 1.03 64.47 7.16 80.15 1.85 72.05 4.01
-RandConv (Xu et al. 2021b) 63.75 4.70 79.35 1.13 65.61 5.07 81.52 1.46 72.56 3.09
-CIDA (Ouyang et al. 2022) 53.16 11.34 78.55 5.18 51.27 14.61 68.09 6.62 62.76 8.44
-PRandConv (Choi et al. 2023) 44.60 9.84 75.23 1.48 53.34 8.68 76.61 2.57 62.44 5.64
+U-Net (ronneberger2015u) 54.82 7.07 77.99 1.16 61.19 8.10 79.49 1.89 68.37 4.56
+AmpSwap (yang2020fda) 63.56 4.36 78.16 1.13 66.13 4.45 81.84 1.45 72.41 2.85
+AmpMix (xu2021fourier) 65.08 4.32 77.06 1.09 68.14 4.29 79.50 1.67 72.44 2.84
+FMAug (li2023frequency) 63.79 5.03 78.34 1.04 64.33 7.28 82.17 1.29 72.16 3.66
+MixStyle (zhou2021mixstyle) 63.95 4.72 80.36 1.06 63.97 6.04 81.93 1.45 72.55 3.32
+TriD (chen2023treasure) 66.27 4.34 77.08 1.24 65.42 4.53 79.36 1.57 72.03 2.92
+ConStyX (chen2025constyx) 62.91 5.99 80.69 1.03 64.47 7.16 80.15 1.85 72.05 4.01
+RandConv (xu2021robust) 63.75 4.70 79.35 1.13 65.61 5.07 81.52 1.46 72.56 3.09
+CIDA (ouyang2022causality) 53.16 11.34 78.55 5.18 51.27 14.61 68.09 6.62 62.76 8.44
+PRandConv (choi2023progressive) 44.60 9.84 75.23 1.48 53.34 8.68 76.61 2.57 62.44 5.64
 StyCona (ours) 67.80 3.60 78.43 1.08 65.71 4.37 80.68 1.49 73.16 2.63
-Figure 3: Qualitative results on the cardiac MRI segmentation and the OC/OD fundus image segmentation tasks.
-Results on Cardiac MRI
-As reported in Table 1, StyCona achieves consistent performance improvements over the compared methods in the cardiac MRI segmentation task.
-In general, almost all the compared methods surpass the baseline model, indicating their effectiveness in addressing domain shifts.
-However, it is worth noting that some augmentation methods, e.g., AmpMix (Xu et al.
-2021a) and TriD (Chen et al.
-2023), yield segmentation results inferior to those of the baseline model in the bSSFP $\rightarrow$ LGE scenario.
-This may stem from the fact that the blended amplitude spectrum in AmpMix and the mixed feature statistics in TriD generate style-augmented images that are insufficient for mitigating content shifts, resulting in the models overfitting spurious correlations between the augmented images and their corresponding segmentation labels.
-In contrast, StyCona demonstrates a clear advantage in handling both style and content shifts by style content augmentation.
-These results validate the core idea of StyCona and highlight its effectiveness in domain-generalizable medical image segmentation.
-Results on OC/OD Fundus Image
-In this experiment, we further evaluate StyCona in a single-source, multi-target domain setting.
-As shown in Table 2, all methods generally achieve comparable average performance across the four target domains.
-However, the two random convolution-based approaches (i.e., CIDA (Ouyang et al.
-2022) and PRandConv (Choi et al.
-2023)) yield unsatisfactory results.
-This may be attributed to that the random convolution operations produce augmented images with distorted content that is inconsistent with the corresponding segmentation labels, thereby misleading the segmentation models during training.
-In contrast, StyCona demonstrates more robust performance across the four target domains, achieving the highest average DSC and competitive ASD.
-These results further suggest the effectiveness of StyCona in domain-generalizable medical image segmentation with multiple target domains.
-Qualitative results.
+3.1.3 Qualitative results.
 As shown in Fig.
@@
 The improvement can be attributed to StyCona’s style and content augmentations, which enable the augmented samples to cover a wide range of unseen domains, thereby allowing the trained model to generate accurate segmentation maps for unseen data.
-In comparison, methods such as AM (Xu et al.
-2021a), MixStyle (Zhou et al.
-2021), and PRandConv (Choi et al.
-2023) yield relatively unsatisfactory results.
-AM and MixStyle rely solely on style augmentation by blending amplitude spectra and adjusting the mean and standard deviation of samples, respectively.
+In comparison, methods such as AmpMix (xu2021fourier), MixStyle (zhou2021mixstyle), and PRandConv (choi2023progressive) yield relatively unsatisfactory results.
+AmpMix and MixStyle rely solely on style augmentation by blending amplitude spectra and adjusting the mean and standard deviation of samples, respectively.
 Meanwhile, the effectiveness of PRandConv is limited by its progressively applied random convolutions, which generate content-distorted images and fail to adequately alleviate variations in anatomical structures.
 These qualitative results, consistent with the quantitative performance, further suggest the effectiveness of our approach.
-Table 3: Ablation study of the proposed method on the cross-sequence setting. The best results are highlighted in bold.
+Figure 3: Qualitative results on the cardiac MRI segmentation and the OC/OD fundus image segmentation tasks.
+3.2 Ablation Study
+3.2.1 Effectiveness of each component
+In Table 3, we present ablation experiments to analyze the contributions of each component of StyCona.
+These experiments were conducted under the cardiac MRI segmentation task.
+The results reveal a consistent trend: segmentation performance improves as the proposed style and content augmentation components are gradually integrated into our method.
+Specifically, the baseline model (i.e., U-Net) performs relatively unsatisfactorily on the target domain, as it tends to learn domain-specific decision rules based on source domain data.
+Then, the segmentation performance increases by a large gap when the style augmentation operation is introduced for effectively mitigating style shifts, which are an essential part of domain shifts in these scenarios.
+In contrast, the improvement is more significant, with an increase of over 7% in DSC, when only the content augmentation component is introduced to address the content shifts, another key aspect of domain shifts.
+Finally, by incorporating both the style and content augmentation modules to address global variations in image appearance and local differences in anatomical structures, our final method achieves the best performance with DSC of 73.39% and 81.59% on the bSSFP $\rightarrow$ LGE and LGE $\rightarrow$ bSSFP scenarios, respectively.
+This result demonstrates the effectiveness of each component and validates our assumptions of the style content decomposition and the corresponding style and content augmentation to alleviate the domain shifts.
+3.2.2 Impact of No. Perturbed Content Maps
+According to Eq.
+(1)(2)(3), perturbing the style codes alters only an image’s style without changing its semantic information corresponding to segmentation labels.
+On the other hand, perturbing $t$ randomly selected content maps results in changes in local anatomical structures (e.g., background tissues or target object-related boundaries), which is the goal of StyCona to enhance the robustness of segmentation models to local anatomical structure variations.
+Qualitatively, Fig.
+4(a) shows the StyCona-augmented samples with varying numbers of perturbed content maps ($t=8,16,32$).
+Compared with the original images, the augmented images with $t=8,16$ exhibit variations in global image appearance and local anatomical structures while remaining consistent with the ground truth segmentation labels.
+In contrast, excessive perturbation is introduced into some target boundaries when $t=32$, which may alter their ground truth labels.
+Quantitatively, $t=16$ yields the highest DSC in both the bSSFP $\rightarrow$ LGE and LGE $\rightarrow$ bSSFP scenarios.
+Based on these results, we set $t=16$ in StyCona to balance perturbation strength with the invariance of segmentation labels, ensuring label consistency after content augmentation.
+Table 3: Ablation study of StyCona on the cross-sequence setting. The best results are highlighted in bold.
 Method StyCona bSSFP $\rightarrow$ LGE LGE $\rightarrow$ bSSFP
@@
 Figure 4: Illustration of a) StyCona-augmented images with different numbers of perturbed content maps ($t=8$, $t=16$, and $t=32$) for two randomly selected original images (The orange dashed circles highlight some content-perturbed regions) and b) Segmentation results for $t=8$, $t=16$, and $t=32$ on cardiac MRI segmentation.
-Ablation Study
-Effectiveness of each component
-In Table 3, we present ablation experiments to analyze the contributions of each component of StyCona.
-These experiments were conducted under the cardiac MRI segmentation task.
-The results reveal a consistent trend: segmentation performance improves as the proposed style and content augmentation components are progressively integrated into our method.
-Specifically, the baseline model (i.e., U-Net) performs relatively unsatisfactorily on the target domain, as it tends to learn domain-specific decision rules based on source domain data.
-Then, the segmentation performance increases by a large gap when the style augmentation operation is introduced, effectively mitigating style shifts, which are an essential part of domain shifts in these scenarios.
-In contrast, the improvement is more significant, with an increase of over 7% in DSC, when only the content augmentation component is introduced to address the content shifts, another key aspect of domain shifts.
-Finally, by incorporating both the style and content augmentation modules to address global variations in image appearance and local differences in anatomical structures, our final method achieves the best performance with DSC of 73.39% and 81.59% on the bSSFP $\rightarrow$ LGE and LGE $\rightarrow$ bSSFP scenarios, respectively.
-This result demonstrates the effectiveness of each component and validates our assumption that style and content augmentations can alleviate the two types of domain shifts.
-Impact of No. Perturbed Content Maps
-According to Eq.
-(1)(2)(3), perturbing the style codes alters only an image’s style without affecting its target shape corresponding to segmentation labels.
-On the other hand, perturbing $t$ randomly selected content maps results in changes in local anatomical structures (e.g., background tissues or target object-related boundaries), which is the goal of our method to enhance the robustness of segmentation models to anatomical structure variations.
-Qualitatively, Fig.
-4(a) shows the StyCona-augmented samples with varying numbers of perturbed content maps ($t=8,16,32$).
-Compared with the original images, the augmented images with $t=8,16$ exhibit variations in overall image appearance and regional anatomical structures while remaining consistent with the ground truth segmentation labels.
-In contrast, at $t=32$, excessive perturbation is introduced into some target boundaries, which may alter their ground truth labels.
-Quantitatively, among all choices, $t=16$ yields the highest DSC in both the bSSFP $\rightarrow$ LGE and LGE $\rightarrow$ bSSFP scenarios.
-Based on these results, we set $t=16$ in StyCona to balance perturbation strength with the invariance of segmentation labels, ensuring label consistency after content augmentation.
-Conclusion
+4 Conclusion
 We propose StyCona, a novel data augmentation algorithm for domain generalizable medical image segmentation.
-Our findings reveal that an image’s singular values function as style codes, governing its global image properties, while its rank-one matrices serve as content maps, determining its anatomical structures.
+Our style content decomposition strategy reveals that an image’s singular values function as style codes, governing its global image properties, while its rank-one matrices serve as content maps, determining its anatomical structures.
 Based on this decomposition, we categorize domain shifts into style and content shifts w.r.t.
 deviations in style codes and content maps, respectively.
-To address these shifts, StyCona perturbs style codes for style augmentation and blends content maps for content augmentation.
-Extensive experiments on cardiac MRI segmentation with a single target domain and OC/OD fundus image segmentation with multiple target domains demonstrate that StyCona is a promising data augmentation technique for domain-generalizable medical image segmentation.
+Then, StyCona perturbs style codes for style augmentation and blends content maps for content augmentation, addressing both types of domain shifts accordingly.
+Extensive experiments on cardiac MRI segmentation with a single target domain and fundus image segmentation with multiple target domains demonstrate that StyCona is an effective data augmentation technique for domain-generalizable medical image segmentation.